# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v2

@@ Abstract
I Introduction
N-Version programming is a software engineering approach to reliability based on crafting software diversity with different teams [1]. It is predominantly used to enhance the reliability of mission-critical systems [2]. This is achieved by building diverse and equivalent software components, with the objective that these would not fail all at once, under the same conditions, hence making the whole system more robust [2].
The largest challenge of N-Version programming relates to the development and operation costs of the approach [3]. There are high costs induced by the manual development and verification of multiple program versions instead of a single one, typically linear in the number of versions. Another important additional cost is the development of additional software to orchestrate the multiple versions (the N-version harness).
Large Language Models (LLMs) are powerful tools able to carry out complex source code-related tasks, such as code completion [4], automatic program repair [5], and code summarization [6]. Our key intuition is that they are able to efficiently automate N-Version programming. Given a reference program, we can leverage the creative and search power of LLMs to automatically produce program variants. We anticipate these variants to be diverse, due to the great volume of data used in LLM training, apt to being combined in an N-Version system. In other words, we want to harness the LLM’s ability to synthesize program variants, and instead of having multiple teams producing multiple versions at a high costs, having the LLM producing them at virtually no cost.
In this paper, we design, implement and evaluate Galápagos, a tool for automated, verified N-Version programming. Galápagos is made up of three passes, which use off-the-shelf LLMs and harness program equivalence verification to provide strong guarantees. Galápagos generates variants of a reference function and then assembles N-Version implementations of that function. We implement Galápagos in the scope of C libraries, with support for function variant generation as well as N-Version in same-language (C-to-C), and cross-language (C-to-Go).
@@ I Introduction
Figure 1: C code which triggers a miscompilation on clang v17.0.0, when using the -Os flag. After running the compiled binary, it prints -1. The correct result is -3. Galápagos is able to detect and mitigate this miscompilation fault.
We conduct a systematic evaluation of Galápagos to quantify the correctness, diversity, and usefulness of both the function variants and the N-Version components. First, we assemble a dataset of 30 real-world C functions, extracted from 6 notable cryptography and multimedia open source projects. We run Galápagos to transform each function in the dataset into an N-Version component. Second, we measure the diversity of the variants, both statically and dynamically. Statically, we measure the uniqueness of the resulting machine code after compiling the variants, using different optimization configurations. Dynamically, we execute each unique variant, and record the CPU instructions they utilize. This allows us to quantify the variation in execution paths, which is known to be important against certain types of exploits such as side-channel attacks [7]. Third, we rigorously test the assembled N-Version implementations in the context of miscompilation bugs [8]. It clearly demonstrates Galápagos’s ability to enhance robustness wrt. our target fault model.
Our experiments demonstrate several key findings: First, Galápagos uses LLMs to generate code variants that are provably equivalent, even when these variants are written in different programming languages. During our experiments, 234 equivalent variants were found from a dataset of 30 reference programs. Second, Galápagos is capable of producing and identifying code variants that exhibit diversity both in their code (statically) and during execution (dynamically). From the dataset of 30 programs, 126 unique variants were verified as different, even after all compilation and optimization passes. Last but not least, these diverse code variants can be utilized to strengthen critical sections of software via automated N-Version programming: we demonstrate Galápagos’ capability to mitigate real-world, reported miscompilation bugs in the Clang compiler.
@@
  • A large scale suite of experiments to evaluate the correctness, value, and practicality of automated N-Version programming with Galápagos. The experiments consist of transforming real-world C code from notable open-source libraries into an N-Version counterpart.
@@ II-A N-Version Programming
N-Version programming is a software development approach, which consists of creating $N$ implementations or versions of a specific program [1]. The core idea behind this approach is that when these versions are simultaneously executed, errors can be timely detected and mitigated by comparing their outputs. Ideally, the difference in implementations between versions is maximal, such that any coincidental errors are avoided. [9]. While originally devised as a fault-tolerance mechanism, N-Version programming has been adapted to enhance other specific properties of software, such as availability [10], reliability [11], performance [12], or security [13, 14]
@@
However, the enhancements offered by N-Version programming come with an attached trade-off, as it introduces many challenges throughout the software development lifecycle. These challenges include increased maintenance overhead, increased compute and memory use, or interoperability issues [15, 10]. Addressing these challenges requires additional effort and careful coordination across engineering teams.
Figure 2: Galapagos pipeline for automated, verified N-Version programming.
@@
An essential challenge is the increased cost of development, given that the time and resources needed to develop the versions increase at least linearly with N. To address this challenge, automating the process for creating new versions is a known and well-studied approach [16, 7, 17, 18]. In this paper, we contribute to this field of automatic synthesis of diverse program versions.
II-B Code Translation with LLMs
@@
Neural machine translation (NMT) for source code refers to the application of machine learning techniques for the automated translation between programming languages [19]. It aims to overcome the challenges associated with traditional rule-based or statistical translation methods by leveraging the power of neural networks to capture and model complex patterns inherent in source code [20]. The core principle behind NMT for source code involves training a neural network model on large datasets consisting of source code snippets in different programming languages and their corresponding translations [21]. During the training phase, the model learns to encode the source code syntax and semantics into a continuous representation, enabling it to generate accurate translations. Subsequently, the trained model can be deployed to translate code snippets from one programming language to another automatically. NMT is tied to recent developments in large language models, which have proven to be efficient in performing translation tasks [22].
III N-Version programming with LLMs
@@
In this section, we present Galápagos, a tool that leverages large language models to harden software against miscompilation errors via automated N-Version programming . Galápagos works at the source-code level for generation, and at the intermediate representation (IR) level for verification. At the source-code level, it uses an LLM to produce new, diversified variants of program functions. At the intermediate representation level, these variants are verified to be semantically equivalent and to be finally assembled into N-Version functions. Because the variants are formally verified to be equivalent, Galápagos is able to provide strong guarantees about the behavior final N-Version assembly, irrespective of the introduced software diversity.
III-A Fault Model
@@
In this paper, we propose a novel technique to mitigate the class of faults called miscompilation. Miscompilations are caused by bugs in compilers, resulting in output machine code not matching the behavior defined in the source code. These faults are very hard to identify, since they are silent, and generally can only be detected indirectly after analyzing misbehaving applications [23].
@@
Miscompilations are present in all major compilers [24]. As of July 2024, Clang and GCC have 340 and 47 unresolved reports of miscompilation bugs, respectively. As an example, consider the code from an open LLVM bug report in Figure 1. After compilation using clang v17.0.0, with the -Os optimization flag, the result is a binary which produces incorrect output: -1. The compilation process exits “successfully” with code 0, and offers no warnings or error messages which would hint that something has gone wrong. Hence, there is an error in one of Clang’s optimization passes. This kind of fault can cause major downstream failures [25].
@@
$P=B_{a}(M(F_{l}(S)))$ (1)
In Equation 1, $S$ is a program’s source code; $F_{l}$ is the front-end stage, which is specific to the language $l$, and transforms $S$ to an intermediate representation (IR); $M$ is the middle-end phase, i.e. a set of optimization passes performed on the IR; $B_{a}$ is the back-end stage, which transforms the IR into assembly code targeted towards a specific architecture ${a}$; and $P$ is the resulting executable program.
@@ III-B Overview
To create an N-Version implementation of the reference function, it is first processed in the Diversification pass (subsection III-C) of Galápagos. Here, an LLM is prompted to automatically generate different variants of the reference. Second, the collection of generated variants is passed to the Validation pass (subsection III-D), which filters out non-viable variants produced by the Diversification pass. The Validation pass proceeds through a sequence of validation steps: (1) Galápagos compiles each variant and produces its corresponding IR code, variants that cannot be compiled are filtered out as non-viable; (2) The IR code of all variants that compile is tested, and sent to a formal equivalence checker to compare their functionality against the reference. The variants that pass the tests and the equivalence check are then forwarded to the next pass of Galápagos. This ensures that after this pass, the ensuing variants have formal guarantees that the original functionality is preserved, and expected diverse internal behavior per the LLM instructions. Third, the Harnessing pass (subsection III-E) uses the resulting variants to assemble an executable, where the original function is replaced by an N-Version implementation of that function. The specific LLM and equivalence checker are external and configurable, which allows Galápagos to use any off-the-shelf solutions.
Galápagos generates a binary that contains an N-Version implementation of the reference function. This binary protects against miscompilation faults, thanks to two properties: (1) miscompilations are very sensitive to the input source code, e.g. they might be related to a specific keyword [26]; consequently, deriving program variants from diverse-but-equivalent source code is less likely to trigger miscompilations in every diversified instance, and; (2) miscompilations are compiler-specific [8], meaning that it is improbable that variants produced by different compilers (see subsection III-C) share the same miscompilation error.
III-C Diversification Pass
@@
The <NUMBER>, <INPUT>, and <OUTPUT> strings are replaced according to the configuration, and the text in square brackets is omitted if the input and output languages are the same.
Figure 3: Galápagos’ diversification pass example. On the left box, the Galápagos prompt; on the right box, the response from the LLM.
Figure 3 shows an example of the diversification pass, including the resulting prompt after configuration, and resulting function variants. In this example, the configuration is set to generate three Go variants from a C reference function. In the response, we can observe that while the first variant is a direct translation of the reference function, the second and third variants are dissimilar, using extra operations or different control flow constructs.
@@
This diversification approach presents two main challenges. First, LLM outputs are not always correct [27]. We need a mechanism to guarantee that the produced variants are equivalent. This concern is addressed in the Validation pass (subsection III-D) Second, when diversifying the reference to different programming languages, the IR code resulting from compiling the variants might not be compatible. We mitigate this concern in the Harnessing pass (subsection III-E).
@@ III-D Validation Pass
In contrast to previous work where automatic diversification is achieved by performing transformations known to preserve equivalence  [26, 16, 7], Galápagos leverages LLMs, and thus the variants are not guaranteed to be correct by construction [28, 29]. To address this issue, Galápagos performs four checks on each variant: compilation in isolation, compilation within containing project, test suite success, and equivalence check.
@@
Equivalence check: In this final check, Galápagos calls a configurable off-the-shelf formal equivalence checking tool, such as alive2 [30], or Rust’s Kani [31]. This uncouples Galápagos from any specific IR language and allows the use of any state-of-the-art solution for verification. It performs the equivalence check with the reference and variant IR codes as parameters. The variant is filtered out if the tool proves the variant to be non-equal, or if it fails to prove equivalence within a given time and memory budget. Reasons for failure at this step are direct evidence of non-equivalence, as the variant deviates from the specification’s functionality. This is the most resource-consuming check, so it is executed last. All function variants that pass the equivalence check are considered correct and can be forwarded to the next pass for assembly. Addressing the accuracy of the equivalence checking tool is out of the scope of this work. We point that equivalence check false negatives would not introduce errors to the system, as the variant would be filtered out. Equivalence check false positives are more challenging, as non-equivalent variants would be accepted in the N-version assembly, Yet, their divergent output would not propagate silent corruptions, but instead would trigger a program termination (see  subsection III-E).
@@ III-E Harnessing Pass
Figure 4: Galápagos’s harnessing pass example, in LLVM IR.
For N-Version programs, the output from the N functions can be selected in different ways. In Galápagos, we support N-of-N selection [32], i.e. we compare the return values of each variant against each other and return a value only if all the variants agree, otherwise Galápagos forcefully terminates the execution.
III-F Class of Functions Considered for Automated N-Version
Galápagos transforms functions into N-Version functions, with equivalence guarantees. To achieve this challenging task, we must assume a specific class of functions. Namely, these functions must be pure since all versions must be executed simultaneously, per the practice of N-Version programming. Only with pure functions can we ensure that multiple executions of the function will not result in accumulated, undesired side effects.
This scope for Galápagos is relevant, as pure functions represent a significant subset of all functions in software [33, 34]. Figure 5 shows a relevant example from the cryptography library libsodium that can be divsersified. Furthermore, we argue that in practice, critical sections of code that have strong reliability requirements can be refactored into pure functions, making hardening via Galápagos and N-Version programming possible.

### v3

# Galápagos: Automated N-Version Programming with LLMs

@@ Abstract.
N-Version Programming, Large Language Models.
$c$s: Software and its engineering Software development techniques
1. Introduction
N-Version programming is a software engineering approach to reliability based on crafting software diversity (building, ) with different teams (avizienis1985n, ). It is predominantly used to enhance the reliability of mission-critical systems (avizienis-2, ). This is achieved by building diverse and equivalent software components, with the objective that these would not fail all at once, under the same conditions, hence making the whole system more robust (avizienis-2, ).
However, N-Version programming faces two fundamental challenges: First, empirical studies have shown that implementations developed by different teams might not be truly independent with respect to the target fault model, as developers tend to share similar mental models and processes, leading to correlated failures (knight2012experimental, ; brilliant2002analysis, ). Second, and the focus of this work, is the significant cost overhead of developing multiple versions. There are high costs induced by the manual development and verification of multiple program versions instead of a single one, typically linear in the number of versions (simplicity, ). Another important additional cost is the development of additional software to orchestrate the multiple versions (the N-Version harness). Our core insight is that LLMs can bring down the cost of engineering N-Version software.
Large Language Models (LLMs) are powerful tools able to carry out complex source code-related tasks, such as code completion (codex, ), automatic program repair (cigar, ), and code summarization (ahmed2022few, ). Our key intuition is that they are able to efficiently automate N-Version programming. Given a reference program, we can leverage the creative and search power of LLMs to automatically produce program variants. We anticipate these variants to be diverse, due to the great volume of data used in LLM training, apt to being combined in an N-Version system. In other words, we want to harness the LLM’s ability to synthesize program variants, and instead of having multiple teams producing multiple versions at a high costs, having the LLM producing them at virtually no cost.
In this paper, we design, implement and evaluate Galápagos, a tool for automated, verified N-Version programming. Galápagos performs three passes: incl. LLM prompting and program equivalence verification to provide strong guarantees. Galápagos generates variants of a reference function and then assembles N-Version implementations of that function. We implement Galápagos to harden C libraries, with support for producing same-language variants (C-to-C), and cross-language variants (C-to-Go).
@@ 1. Introduction
Figure 1. C code which triggers a miscompilation on clang v17.0.0, when using the -Os flag (left). After running the compiled binary, it prints -1. The correct result is -3 (right). Galápagos is able to detect and mitigate this miscompilation fault.
We conduct a systematic evaluation of Galápagos to quantify the correctness, diversity, and usefulness of both the function variants and the N-Version components. First, we assemble a dataset of 30 real-world C functions, extracted from 6 notable cryptography and multimedia open source projects. We run Galápagos to transform each function in the dataset into an N-Version component. Second, we measure the diversity of the variants, both statically and dynamically. Statically, we measure the uniqueness of the resulting machine code after compiling the variants, using different optimization configurations. Dynamically, we execute each unique variant, and record the CPU instructions they utilize, comparing against state-of-the-art diversification and obfuscation tool Tigress (tigress, ) as a baseline. This allows us to quantify the variation in execution paths, which is known to be important against certain types of exploits such as side-channel attacks (wasm-mutate, ). Third, we rigorously test the assembled N-Version implementations in the context of miscompilation bugs (yang2011finding, ). It clearly demonstrates Galápagos’s ability to enhance robustness wrt. our target fault model.
Our experiments demonstrate several key findings: First, Galápagos uses LLMs to generate code variants that are provably equivalent, even when these variants are written in different programming languages. During our experiments, 234 equivalent variants were found from a dataset of 30 reference programs. Second, Galápagos is capable of producing and identifying code variants that exhibit diversity both in their code (statically) and during execution (dynamically), outperforming transformation-based diversification techniques like Tigress. From the dataset of 30 programs, 126 unique variants were verified as different, even after all compilation and optimization passes. Last but not least, these diverse code variants can be utilized to strengthen critical sections of software via automated N-Version programming: we demonstrate Galápagos’ capability to mitigate real-world, reported miscompilation bugs in the Clang compiler.
@@
  • A systematic suite of experiments to evaluate the correctness, value, and practicality of automated N-Version programming with Galápagos. The experiments consist of transforming real-world C code from notable open-source libraries into an N-Version counterpart.
@@ 2.1. N-Version Programming
N-Version programming is a software development approach, which consists of creating $N$ implementations or versions of a specific program (avizienis1985n, ). The core idea behind this approach is that when these versions are simultaneously executed, errors can be timely detected and mitigated by comparing their outputs. Ideally, the difference in implementations between versions is maximal, such that any coincidental errors are avoided. (1701974, ). While originally devised as a fault-tolerance mechanism, N-Version programming has been adapted to enhance other specific properties of software, such as availability (neth, ), reliability (error-recovery, ; verifiable-client-diversity, ), performance (n-version-perf, ), or security (n-version-sec, ; back-to-the-future, )
@@
However, the enhancements offered by N-Version programming come with an attached trade-off, as it introduces many challenges throughout the software development lifecycle. These challenges include increased maintenance overhead, increased compute and memory use, or interoperability issues (n-version-challenges, ; neth, ). Addressing these challenges requires additional effort and careful coordination across engineering teams.
Figure 2. Galápagos pipeline for automated, verified N-Version programming.
@@
An essential challenge is the increased cost of development, given that the time and resources needed to develop the versions increase at least linearly with N. To address this challenge, automating the process for creating new versions is a known and well-studied approach (cadar-nvx, ; wasm-mutate, ; n-version-obf, ; los-alamos, ). In this paper, we contribute to this field of automatic synthesis of diverse program versions.
2.2. Code Translation with LLMs
@@
Neural machine translation (NMT) for source code refers to the application of machine learning techniques for the automated translation between programming languages (tufano-nmt, ). It aims to overcome the challenges associated with traditional rule-based or statistical translation methods by leveraging the power of neural networks to capture and model complex patterns inherent in source code (sequencer, ). The core principle behind NMT for source code involves training a neural network model on large datasets consisting of source code snippets in different programming languages and their corresponding translations (facebook-nmt, ). During the training phase, the model learns to encode the source code syntax and semantics into a continuous representation, enabling it to generate accurate translations. Subsequently, the trained model can be deployed to translate code snippets from one programming language to another automatically. NMT is tied to recent developments in large language models, which have proven to be efficient in performing translation tasks (vert, ).
3. N-Version programming with LLMs
@@
In this section, we present Galápagos, a tool that leverages large language models to harden software via automated N-Version programming. Galápagos works at the source-code level for generation, and at the intermediate representation (IR) level for verification. At the source-code level, it uses an LLM to produce new, diversified variants of program functions. At the intermediate representation level, these variants are verified to be semantically equivalent and to be finally assembled into N-Version functions. Because the variants are formally verified to be equivalent, Galápagos is able to provide strong guarantees about the behavior final N-Version assembly, irrespective of the introduced software diversity.
3.1. Fault Model
@@
In this paper, we propose a novel technique to mitigate the class of faults called miscompilation. Miscompilations occur when a compiler erroneously generates machine code that deviates from the behavior defined in the source code, due to internal bugs in the compiler itself. These compiler faults are very hard to identify because they are typically silent. They are critical because the resulting application binaries behave incorrectly at runtime (sun2016toward, ).
@@
Miscompilations are present in all major compilers (compcert, ). As of May 2025, Clang and GCC have 396 and 44 unresolved reports of miscompilation bugs, respectively. As an example, consider the code from an open LLVM bug report in Figure 1. After compilation using clang v17.0.0, with the -Os optimization flag, the result is a binary which produces incorrect output: -1. The compilation process exits “successfully” with code 0, and offers no warnings or error messages which would hint that something has gone wrong. Hence, there is an error in one of Clang’s optimization passes. This kind of fault can cause major downstream failures (marcozzi2019compiler, ).
@@
(1) $P=B_{a}(M(F_{l}(S)))$
In Equation 1, $S$ is a program’s source code; $F_{l}$ is the front-end stage, which is specific to the language $l$, and transforms $S$ to an intermediate representation (IR); $M$ is the middle-end phase, i.e. a set of optimization passes performed on the IR; $B_{a}$ is the back-end stage, which transforms the IR into assembly code targeted towards a specific architecture a{a}; and $P$ is the resulting executable program. Furthermore, $M$ is subject to a wide range of configuration options, each of which can potentially introduce miscompilation bugs. For example, optimization flags such as -Os or -O3 enable aggressive transformations intended to reduce code size or improve performance, but they also increase the risk of activating miscompilation bugs in these optimization passes.
@@ 3.2. Overview
To create an N-Version implementation of the reference function, it is first processed in the Diversification pass (subsection 3.3) of Galápagos. Here, an LLM is prompted to automatically generate different variants of the reference. Second, the collection of generated variants is passed to the Validation pass (subsection 3.4), which filters out non-viable variants produced by the Diversification pass. The Validation pass proceeds through a sequence of validation steps: (1) Galápagos compiles each variant and produces its corresponding IR code, variants that cannot be compiled are filtered out as non-viable; (2) The IR code of all variants that compile is tested, and sent to a formal equivalence checker to compare their functionality against the reference. The variants that pass the tests and the equivalence check are then forwarded to the next pass of Galápagos. This ensures that after this pass, the ensuing variants have formal guarantees that the original functionality is preserved, and expected diverse internal behavior per the LLM instructions. Third, the Harnessing pass (subsection 3.5) uses the resulting variants to assemble an executable, where the original function is replaced by an N-Version implementation of that function. The specific LLM and equivalence checker are external and configurable, which allows Galápagos to use any off-the-shelf solutions.
Galápagos generates a binary that contains an N-Version implementation of the reference function. This binary protects against miscompilation faults, thanks to two properties: (1) miscompilations are very sensitive to the input source code, e.g. they might be related to a specific keyword (regehr-volatile, ); consequently, deriving program variants from diverse-but-equivalent source code is less likely to trigger miscompilations in every diversified instance, and; (2) miscompilations are compiler-specific (yang2011finding, ), meaning that it is improbable that variants produced by different compilers (see subsection 3.3) share the same miscompilation error.
3.3. Diversification Pass
@@
The <NUMBER>, <INPUT>, and <OUTPUT> strings are replaced according to the configuration, and the text in square brackets is omitted if the input and output languages are the same. Note that some LLM APIs provide support for sampling n times with the same prompt (aka n_sample). Our pilot experiment showed that a single prompt with n_sample > 1 produced less diverse results than one single prompts requesting for <NUMBER> variant. Hence, this is what we use in this paper.
Figure 3. Galápagos’ diversification pass example. On the left box, the Galápagos prompt; on the right box, the response from the LLM.
Figure 3 shows an example of the diversification pass, including the resulting prompt after configuration, and resulting function variants. The configuration is set to generate three Go variants from a C reference function. Recall that translating to another language such as Go, provides runtime diversity which is the core goal of N-Version programming (avizienis1985n, ).
In the response, we can observe that while the first variant is a direct translation of the reference function, the second and third variants are dissimilar, using extra operations or different control flow constructs.
@@
This diversification approach presents two main challenges. First, LLM outputs are not always correct (creative-correct, ). We need a mechanism to guarantee that the produced variants are equivalent. This concern is addressed in the Validation pass (subsection 3.4) Second, when diversifying the reference to different programming languages, the IR code resulting from compiling the variants might not be compatible. We mitigate this concern in the Harnessing pass (subsection 3.5).
@@ 3.4. Validation Pass
In contrast to previous work where automatic diversification is achieved by performing transformations (1) that operate on low-level code representations, e.g. binary code, and (2) that are known to preserve equivalence (nop, ; kguard, ; wasm-mutate, ), Galápagos leverages LLMs at the source code-level, and thus the variants are not guaranteed to be correct by construction (NEURIPS2023_43e9d647, ; liu2024exploring, ). To address this issue, Galápagos performs four checks on each variant: compilation in isolation, compilation within containing project, test suite success, and equivalence check.
@@
Equivalence check: In this final check, Galápagos calls a configurable off-the-shelf formal equivalence checking tool, such as alive-tv (lopes2021alive2, ), or Rust’s Kani (kani, ). This uncouples Galápagos from any specific IR language and allows the use of any state-of-the-art solution for verification. It performs the equivalence check with the reference and variant IR codes as parameters. The variant is filtered out if the tool proves the variant to be non-equal, or if it fails to prove equivalence within a given time and memory budget. Reasons for failure at this step are direct evidence of non-equivalence, as the variant deviates from the specification’s functionality. This is the most resource-consuming check, so it is executed last. All function variants that pass the equivalence check are considered correct and can be forwarded to the next pass for assembly. Addressing the accuracy of the equivalence checking tool is out of the scope of this work. We point that equivalence check false negatives would not introduce errors to the system, as the variant would be filtered out. Equivalence check false positives are more challenging, as non-equivalent variants would be accepted in the N-Version assembly, yet, their divergent output would not propagate silent corruptions, but instead would trigger a program termination (see  subsection 3.5).
@@ 3.5. Harnessing Pass
Figure 4. Galápagos’s harnessing pass example, in LLVM IR.
For N-Version programs, the output from the N functions can be selected in different ways. In Galápagos, we support N-of-N selection (hydra, ), i.e. we compare the return values of each variant against each other and return a value only if all the variants agree, otherwise Galápagos forcefully terminates the execution.
3.6. Class of Functions Considered for Automated N-Version
Galápagos transforms functions into N-Version functions, with equivalence guarantees. To achieve this challenging task, we must assume a specific class of functions. Namely, these functions must be pure, i.e. must deterministically produce the same output from the same input, and must not cause any side-effect. This is a requisite to our N-Version approach since all function versions must be executed simultaneously, and their outputs compared. Only with pure functions can we ensure that multiple executions of the function will not result in the accumulation of undesired side effects.
This scope for Galápagos is relevant, as pure functions represent a significant subset of all functions in software (pure-java, ; pure-javascript, ). Figure 5 shows a relevant example from the cryptography library libsodium that can be divsersified. Furthermore, we argue that in practice, critical sections of code that have strong reliability requirements can be refactored into pure functions, making hardening via Galápagos and N-Version programming possible.